OpenStudy (anonymous):

The perimeter of a triangle is 11y - 2. Two of the sides are represented by the expression 3y - 1 and 3y + 1. Write an expression for the third side.

OpenStudy (texaschic101):

perimeter of a triangle is : P = a + b + c P = 11y - 2 a = 3y - 1 b = 3y + 1 c = 3rd side 11y - 2 = (3y - 1) + (3y + 1) + c 11y - 2 = 6y + c (11y - 2)/6y = c I think ...you might want to get a second opinion

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (texaschic101):

or 11y - 2 = 6y + c 11y - 6y - 2 = c 7y - 2 = c

OpenStudy (anonymous):

These are the choices a. 6y c. 5y - 2 b. 6y - 2 d. -2y

OpenStudy (texaschic101):

oops...I subtracted wrong 11y - 2 = 6y + c 11y - 6y - 2 = c 5y - 2 = c

OpenStudy (texaschic101):

I messes up on the first explanation....just disregard that one

OpenStudy (texaschic101):

typo...messed up

OpenStudy (anonymous):

Thank you got the answer correct you're a life saver!